# Black-hole scattering with numerical relativity: Self-force extraction and post-Minkowskian validation

### v1

@@ I Introduction
The most natural description of small deflection scattering is through the post-Minkowskian (PM) expansion, which treats the gravitational interaction perturbatively in powers of Newton’s constant $G$. The vast majority of recent theoretical efforts in this area have focused on calculating high-order PM contributions to scattering observables using a variety of techniques borrowed from high-energy physics, including scattering amplitudes Bern et al. (2021, 2022); Damgaard et al. (2023) and worldline quantum field theory Dlapa et al. (2022, 2023); Jakobsen et al. (2023, 2024); Driesse et al. (2024a, b). These calculations have provided valuable insights into the dynamics of BBH encounters and have been used to construct bound-orbit waveform models Buonanno et al. (2024a); Damour et al. (2025). However, these calculations break down when approaching the strong field regime, where the curvature of spacetime becomes significant, and non-perturbative effects dominate the dynamics. In order to remain faithful at small separations, one can apply resummation techniques such as effective-one-body (EOB) Damour and Rettegno (2023); Buonanno et al. (2024b, a); Damour et al. (2025); Long et al. (2025); Clark and Pratten (2025) or incorporate information about the scatter-capture separatrix Damour and Rettegno (2023); Long et al. (2024).
@@
The PM and SF frameworks are perturbative by construction, and determining their accuracy requires comparison with fully non-perturbative results. Numerical relativity (NR) provides this benchmark by directly solving discretized versions of the Einstein field equations on high-performance computers. These simulations provide a detailed, fully non-linear description of phenomena and capture all of the physics of the system. While NR codes were originally designed to model the late inspiral and merger of compact binaries, they have recently been adapted to study scattering encounters Damour et al. (2014); Damour and Rettegno (2023); Hopper et al. (2023); Rettegno et al. (2023); Albanesi et al. (2025); Swain et al. (2025); Fontbuté et al. (2025a); Pretorius and Khurana (2007); Sperhake et al. (2008); Witek et al. (2010); Sperhake et al. (2011, 2013, 2016); Kankani and McWilliams (2024); Nelson et al. (2019); Jaraba and Garcia-Bellido (2021); Rodríguez-Monteverde et al. (2025); Bae et al. (2024); Fontbuté et al. (2025b); Long et al. (2025); Kogan et al. (2025).
@@ II Numerical relativity simulations
SpEC was recently adapted to perform accurate simulations of hyperbolic encounters. These changes included modifications to the initial data routines, to account for larger separations and velocities, as well as the adaptive mesh refinement procedure during the evolution. Details of these modifications can be found in Refs. Long et al. (2025); Mendes et al. (2025).
In this study we analyze the observables from 23 simulations previously presented in Ref. Long et al. (2025), combined with 9 new equal-mass simulations at fixed $\Gamma=1.02264$ deeper in the weak field (impact parameters $\ell\geq 8$ corresponding to $b\geq 18.7M$). A subset of the equal-mass simulations are shown in Figure 1. The rest of this section is dedicated to summarizing the details of the simulations performed and our procedure for extracting the scattering angle from the data.
Figure 1: Trajectories of equal-mass black-hole scattering scenarios with $\Gamma=1.02264$ and impact parameters $b=[37.4,30.4,23.4,17.1,12.8,11.1]M$. Shown are the Cartesian components of the separation vector between the centers of the BHs, rotated such that the incoming path is aligned with the $x$-axis. The $b=17.1M$ (red) trajectory corresponds to the largest impact parameter previously studied at this energy Damour et al. (2014); Rettegno et al. (2023); Swain et al. (2025); Long et al. (2025).
@@ II.1 Simulations
The setup of all simulations used here is exactly as detailed in Sec. II of Ref. Long et al. (2025), to summarize:
Each simulation is specified by the parameters of the BHs (masses and spins) combined with parameters that determine the trajectories of the binary, notably the ADM energy and angular momentum, $E_{\rm ADM}$ and $J_{\rm ADM}$, as well as the initial separation $D_{0}=250M$. In order to evolve this system, we solve for initial data on the first slice of the evolution. The SpEC initial data routine Pfeiffer et al. (2003); Ossokine et al. (2015) performs a root-finding procedure over additional parameters to achieve the desired system. This includes those that control the initial position and velocity of the BHs as well as the conformal metric, trace of the extrinsic curvature and their time-derivatives and the boundary conditions at the BH horizons and spatial infinity. As discussed in Ref. Long et al. (2025), we can achieve high accuracy in the desired values of the masses, spins, ADM energy and angular momentum, etc., with relative errors $\lesssim 10^{-5}$, ensuring that the results here are not affected by uncertainty in the initial parameters.
@@ II.2 Scattering angle extraction
The scattering angle $\theta$ is a gauge-invariant quantity that can be used to compare different simulations and analytic predictions. In order to calculate this quantity we must track the motion of each BH throughout the simulation and extrapolate their trajectories to infinite separation. Here we review the extrapolation procedure and refer the reader to Sec. II of Ref. Long et al. (2025) for a detailed explanation.
@@
For the equal mass case considered in Sec. IV, this is equivalent to the individual scattering angles, whereas for the unequal mass case in Sec. III, the broken symmetry of the system and the associated asymmetric gravitational wave emission means that $\theta_{1}\neq\theta_{2}$. However, as shown in Ref. Long et al. (2025), the difference in the angles of each BH is small compared to the overall scattering angle ($\lesssim 0.1\%$) and thus we neglect the difference here.
@@ III Self-force regime
The leading-order term $\vartheta_{\rm 0SF}$ represents the scattering angle of a geodesic in a Schwarzschild spacetime and can be calculated using elliptic integrals Barack and Long (2022). The higher-order terms, $\vartheta_{n\text{SF}}$ for $n\geq 1$, represent corrections due to the finite mass ratio of the binary system. These terms are significantly more challenging to compute and thus have only been calculated with additional approximations, such as the latest PM results which contain information at $G^{5}$ up to linear order in the mass ratio, i.e. 5PM(1SF) Driesse et al. (2024b, a).
@@
where $\vartheta_{n\text{PM}}$ is the $n$th order PM contribution to the scattering angle and we have restored Newton’s constant $G$ for clarity. The state-of-the-art results are at 5PM(1SF) which contain all information up to $G^{4}$ Bern et al. (2021) as well as the contributions up to order $\nu$ at $G^{5}$ Driesse et al. (2024b, a). In addition, as the $\nu^{0}$ contributions are purely geodesic, we have these contributions for all orders in $G$.
@@
Another informative way to visualize the agreement between NR and PM is to plot the NR results with successive PM orders subtracted off, as shown in Figure 5. If there is agreement, each set of points should tend towards the corresponding PM prediction with increasing $b$. As can be seen, the NR results agree well with the PM predictions at each order, providing a strong validation between the NR and PM calculations. One interesting observation is that when subtracting the 5PM(1SF) term, the points appear to follow a trend of $\sim 1/b^{6}$ at large $b$. This suggests that the next order term, 6PM, is larger than the missing 2SF contribution at 5PM. In fact, when comparing these results to the known 6PM $\nu^{0}$ contribution, we find that this is indeed the case and that the 6PM(0SF) term is actually within the error bars of the NR data at large $b$. However, this does not mean that the whole 6PM term is within the error bars as the unknown higher-order in $\nu$ contributions could be significant.
@@ Appendix A Numerical values
In this appendix, we provide the numerical values of the scattering angles from our SpEC simulations used in Secs. III and IV. The ancillary file contains machine-readable versions of these tables with the exact initial conditions and extracted scattering angle values for each simulation.
@@
Table 3: Scattering angle (in degrees) for equal mass, non-spinning BHs with $\Gamma\!=\!1.02264$. $\theta_{n\text{PM}}$ is the analytic PM prediction containing terms up to the at $n$th-order i.e. $\theta_{n\text{PM}}=\sum_{i=1}^{n}\vartheta_{i\text{PM}}\left(\frac{GM}{b}\right)^{i}$. $\theta_{5\text{PM}}^{1\text{SF}}$ represents $\theta_{5\text{PM}}$ truncated to terms up to order $\nu$ as the $\nu^{2}$ terms are currently unknown Driesse et al. (2024b, a).

### v2

@@ I Introduction
The most natural description of small deflection scattering is through the post-Minkowskian (PM) expansion, which treats the gravitational interaction perturbatively in powers of Newton’s constant $G$. The vast majority of recent theoretical efforts in this area have focused on calculating high-order PM contributions to scattering observables using a variety of techniques borrowed from high-energy physics, including scattering amplitudes Bern et al. (2021, 2022); Damgaard et al. (2023) and worldline field theory Kälin and Porto (2020); Dlapa et al. (2022, 2023a, 2023b); Jakobsen et al. (2023, 2024); Driesse et al. (2025, 2024). These calculations have provided valuable insights into the dynamics of BBH encounters and have been used to construct bound-orbit waveform models Buonanno et al. (2024a); Damour et al. (2025). However, these calculations break down when approaching the strong field regime, where the curvature of spacetime becomes significant, and non-perturbative effects dominate the dynamics. In order to remain faithful at small separations, one can apply resummation techniques such as effective-one-body (EOB) Damour and Rettegno (2023); Buonanno et al. (2024b, a); Damour et al. (2025); Long et al. (2025a); Clark and Pratten (2025) or incorporate information about the scatter-capture separatrix Damour and Rettegno (2023); Long et al. (2024).
@@
The PM and SF frameworks are perturbative by construction, and determining their accuracy requires comparison with fully non-perturbative results. Numerical relativity (NR) provides this benchmark by directly solving discretized versions of the Einstein field equations on high-performance computers. These simulations provide a detailed, fully non-linear description of phenomena and capture all of the physics of the system. While NR codes were originally designed to model the late inspiral and merger of compact binaries, they have recently been adapted to study scattering encounters Damour et al. (2014); Damour and Rettegno (2023); Hopper et al. (2023); Rettegno et al. (2023); Albanesi et al. (2025); Swain et al. (2025); Fontbuté et al. (2025a); Pretorius and Khurana (2007); Sperhake et al. (2008); Witek et al. (2010); Sperhake et al. (2011, 2013, 2016); Kankani and McWilliams (2024); Nelson et al. (2019); Jaraba and Garcia-Bellido (2021); Rodríguez-Monteverde et al. (2025); Bae et al. (2024); Fontbuté et al. (2025b); Long et al. (2025a); Kogan et al. (2025).
@@ II Numerical relativity simulations
SpEC was recently adapted to perform accurate simulations of hyperbolic encounters. These changes included modifications to the initial data routines, to account for larger separations and velocities, as well as the adaptive mesh refinement procedure during the evolution. Details of these modifications can be found in Refs. Long et al. (2025a); Mendes et al. (2025).
In this study we analyze the observables from 23 simulations previously presented in Ref. Long et al. (2025a), combined with 9 new equal-mass simulations at fixed $\Gamma=1.02264$ deeper in the weak field (impact parameters $\ell\geq 8$ corresponding to $b\geq 18.7M$). A subset of the equal-mass simulations are shown in Figure 1. The rest of this section is dedicated to summarizing the details of the simulations performed and our procedure for extracting the scattering angle from the data.
Figure 1: Trajectories of equal-mass black-hole scattering scenarios with $\Gamma=1.02264$ and impact parameters $b=[37.4,30.4,23.4,17.1,12.8,11.1]M$. Shown are the Cartesian components of the separation vector between the centers of the BHs, rotated such that the incoming path is aligned with the $x$-axis. The $b=17.1M$ (red) trajectory corresponds to the largest impact parameter previously studied at this energy Damour et al. (2014); Rettegno et al. (2023); Swain et al. (2025); Long et al. (2025a).
@@ II.1 Simulations
The setup of all simulations used here is exactly as detailed in Sec. II of Ref. Long et al. (2025a), to summarize:
Each simulation is specified by the parameters of the BHs (masses and spins) combined with parameters that determine the trajectories of the binary, notably the ADM energy and angular momentum, $E_{\rm ADM}$ and $J_{\rm ADM}$, as well as the initial separation $D_{0}=250M$. In order to evolve this system, we solve for initial data on the first slice of the evolution. The SpEC initial data routine Pfeiffer et al. (2003); Ossokine et al. (2015) performs a root-finding procedure over additional parameters to achieve the desired system. This includes those that control the initial position and velocity of the BHs as well as the conformal metric, trace of the extrinsic curvature and their time-derivatives and the boundary conditions at the BH horizons and spatial infinity. As discussed in Ref. Long et al. (2025a), we can achieve high accuracy in the desired values of the masses, spins, ADM energy and angular momentum, etc., with relative errors $\lesssim 10^{-5}$, ensuring that the results here are not affected by uncertainty in the initial parameters.
@@ II.2 Scattering angle extraction
The scattering angle $\theta$ is a gauge-invariant quantity that can be used to compare different simulations and analytic predictions. In order to calculate this quantity we must track the motion of each BH throughout the simulation and extrapolate their trajectories to infinite separation. Here we review the extrapolation procedure and refer the reader to Sec. II of Ref. Long et al. (2025a) for a detailed explanation.
@@
For the equal mass case considered in Sec. IV, this is equivalent to the individual scattering angles, whereas for the unequal mass case in Sec. III, the broken symmetry of the system and the associated asymmetric gravitational wave emission means that $\theta_{1}\neq\theta_{2}$. However, as shown in Ref. Long et al. (2025a), the difference in the angles of each BH is small compared to the overall scattering angle ($\lesssim 0.1\%$) and thus we neglect the difference here.
@@ III Self-force regime
The leading-order term $\vartheta_{\rm 0SF}$ represents the scattering angle of a geodesic in a Schwarzschild spacetime and can be calculated using elliptic integrals Barack and Long (2022). The higher-order terms, $\vartheta_{n\text{SF}}$ for $n\geq 1$, represent corrections due to the finite mass ratio of the binary system. These terms are significantly more challenging to compute and thus have only been calculated with additional approximations, such as the latest PM results which contain information at $G^{5}$ up to linear order in the mass ratio, i.e. 5PM(1SF) Driesse et al. (2024, 2025).
@@
where $\vartheta_{n\text{PM}}$ is the $n$th order PM contribution to the scattering angle and we have restored Newton’s constant $G$ for clarity. The state-of-the-art results are at 5PM(1SF) which contain all information up to $G^{4}$ Bern et al. (2021); Dlapa et al. (2023a) as well as the contributions up to order $\nu$ at $G^{5}$ Driesse et al. (2024, 2025). In addition, as the $\nu^{0}$ contributions are purely geodesic, we have these contributions for all orders in $G$.
@@
Another informative way to visualize the agreement between NR and PM is to plot the NR results with successive PM orders subtracted off, as shown in Figure 5. If there is agreement, each set of points should tend towards the corresponding PM prediction with increasing $b$. As can be seen, the NR results agree well with the PM predictions at each order, providing a strong validation between the NR and PM calculations. One interesting observation is that when subtracting the 5PM(1SF) term, the points appear to follow a trend of $\sim 1/b^{6}$ at large $b$. This suggests that the next order term, 6PM, is larger than the missing 2SF contribution at 5PM. In fact, when comparing these results to the known 6PM $\nu^{0}$ contribution, we find that this is indeed the case and that the 6PM(0SF) term is actually within the error bars of the NR data at large $b$. While we expect that the $\nu^{0}$ contribution dominates the 6PM term, as has been observed at lower PM orders (see e.g. Fig. 7 of Ref. Long et al. (2025a)), this does not mean that the whole 6PM term is within the error bars as the unknown higher-order in $\nu$ contributions could still be significant.
@@ Acknowledgements
Data Availability
An ancillary file containing the scattering angle data used in this article is openly available at Long et al. (2025b).
@@ Appendix A Numerical values
In this appendix, we provide the numerical values of the scattering angles from our SpEC simulations used in Secs. III and IV. The ancillary file Long et al. (2025b) contains machine-readable versions of these tables with the exact initial conditions and extracted scattering angle values for each simulation.
@@
Table 3: Scattering angle (in degrees) for equal mass, non-spinning BHs with $\Gamma\!=\!1.02264$. $\theta_{n\text{PM}}$ is the analytic PM prediction containing terms up to the at $n$th-order i.e. $\theta_{n\text{PM}}=\sum_{i=1}^{n}\vartheta_{i\text{PM}}\left(\frac{GM}{b}\right)^{i}$. $\theta_{5\text{PM}}^{1\text{SF}}$ represents $\theta_{5\text{PM}}$ truncated to terms up to order $\nu$ as the $\nu^{2}$ terms are currently unknown Driesse et al. (2024, 2025).